The area of a rectangle garden is 36 meters.the width is 4 m longer than one third of the length find the length and the width of the garden
Answer by Cromlix(4381)

Hi there,
Make length = x
Width = 4 + x/3
Area = Length x width
36 = x(4 + x/3)
36 = 4x + x^2/3
Rearrange into y = ax^2 + bx + c
x^2/3 + 4x - 36 = 0
Multiply throughout by 3
x^2 + 12x - 108 = 0
(x + 18)(x - 6) = 0
x + 18 = 0
x = -18 (Disregard as -ve)
x - 6 = 0
x = 6
Length = 6m
Width = 6m
Hope this helps :-)